OpenStudy (anonymous):

2/3 of what number is 16?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

(2/3)x=16 x= 16 (3/2) x= (16 x 3)/2 x= 48/2 x=24

OpenStudy (anonymous):

Welcome to OpenStudy @tlambert23 !

OpenStudy (anonymous):

thanks